OpenStudy (yueyue):

Find the limit of the function algebraically. lim x^2+2x/x^4 x->0

OpenStudy (amistre64):

algebracilly eh, divide top and bottom by x^4

OpenStudy (yueyue):

I've tried it a million times but I always get zero in the denominator. I think my teacher wants me to find one without a zero as the denominator.

OpenStudy (amistre64):

or is this something to do with "first princilpes" of derivatives

OpenStudy (yueyue):

Yes it deals with derivatives.

OpenStudy (amistre64):

.... then you are in for a nightmare lol

OpenStudy (amistre64):

\[\lim_{h\to 0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (amistre64):

i hear that sometimes: f(b)-f(a) ------ is simpler b-a

OpenStudy (yueyue):

Ah I don't know if I must use that expression for this question but I'll give it a try.

OpenStudy (yueyue):

Yeah, definitely not how I am supposed to be doing it hahah. Thank you though @amistre64

OpenStudy (dumbcow):

\[\frac{x^{2}+2x}{x^{4}} = \frac{x(x+2)}{x^{4}} = \frac{x+2}{x^{3}}\] \[\lim_{x \rightarrow 0} \frac{x+2}{x^{3}} = \lim_{x \rightarrow 0} \frac{2}{0}=\infty\]

OpenStudy (yueyue):

Yes okay good I figured there is no way to make the denominator anything other than zero but I wasn't sure. Thank you so much for confirming it! @dumbcow

OpenStudy (dumbcow):

yw

OpenStudy (anonymous):

Doe that mean it is zero, or that it does not exist?

OpenStudy (anonymous):

*Does

OpenStudy (amistre64):

well, infinity is not zero, so it aint zero. so it depends on how the course defines a divide by zero. DNE is fine for some classes, and infinity is fine for others.

OpenStudy (anonymous):

DOES NOT EXIST